OpenStudy (anonymous):

Please help will fan and give a medal What is the value of the expression 7^-3 A. -21 b.-1/343 c.1/343 d.343 btw -3 is a exponet

OpenStudy (anonymous):

\[7^{-3} = \frac{ 1 }{ 7^{3} }\]

OpenStudy (anonymous):

that's a rule when it comes to negative exponents

OpenStudy (anonymous):

ok thank you but im still confused

OpenStudy (anonymous):

what's confusing you ?

OpenStudy (anonymous):

o ok so the answer is d?

OpenStudy (anonymous):

im not exactly sure what im confused about but I just took a quiz on it an mad a 60%

OpenStudy (anonymous):

ok let's say we have 7^ 3 that's 7 x 7 x7 right ?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

okay when you have a negative exponent it's \[\frac{ 1 }{ 7 x7 x7 }\]

OpenStudy (anonymous):

the xs are meant to be times sorry about that

OpenStudy (anonymous):

so the answer is 343 because 7*7*7 is 343

OpenStudy (anonymous):

no that answer isn't 343 your question says 7^-3 meaning 7 to the power of negative 3

OpenStudy (anonymous):

o ok

OpenStudy (anonymous):

if your question was 7 to the power of 3 ( as in positive three) the answer would be 343 let me say it in another way positive exponents : eg. 3^2 = 3 x 3 negative exponents : eg. 3^-2 = \[1\div3\div3\] in other words \[\frac{ 1 }{ 3\times3 }\]

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

yes thank you sooo much for ur help

OpenStudy (anonymous):

got the answer ?

OpenStudy (anonymous):

b I think